OpenStudy (omarbirjas):

Solve each equation in the interval from 0 to 2pi. Round your answer to the nearest hundreth.

OpenStudy (omarbirjas):

\[\cos (\pi ~ \theta) = -0.6\]

OpenStudy (omarbirjas):

@dtan5457